# Beyond-mean-field phases of rotating dipolar condensates

### v1

@@ I Introduction
The recent progress in experiments with dipolar ultracold atoms has provided a platform for studying novel quantum phases emerging from the interplay between contact and long-range dipole-dipole interactions (DDI) [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13]. These systems have enabled the realization of quantum droplets [14, 15] and supersolid phases [16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27, 28, 29], as well as the observation of dipolar quantum gases with metastable properties [30, 31, 32]. These interactions give rise to striking phenomena such as unconventional vortex clustering and dynamics [33, 34, 35, 36], roton instabilities [37], and self-bound quantum droplets [15, 14].
@@
Beyond these static properties, the interplay between dipolar interactions and external rotation introduces even richer physics. Recent studies have demonstrated that rotating dipolar gases exhibit complex vortex arrangements, turbulence, and dynamically stabilized cluster states [38, 39, 40, 41, 42, 43, 44]. A particularly intriguing aspect of rotating dipolar systems is their connection to astrophysical phenomena. In neutron stars, superfluid vortices exhibit long-range interactions that shape the internal structure and rotational properties of the star [45, 46]. Understanding the interplay between rotation and dipolar interactions is therefore essential for advancing both laboratory-based quantum simulations and our knowledge of astrophysical superfluid systems.
@@
The presence of strong dipolar interactions modifies the well-known vortex lattice structures observed in non-dipolar BECs. In contrast to conventional Abrikosov lattices, dipolar condensates can exhibit striped vortex arrangements, fragmented clusters, or even phase-separated vortex distributions [35, 34]. Theoretical studies have predicted exotic vortex lattice phases, including triangular, square, and amorphous configurations, depending on the interaction strength and rotation frequency [41, 44]. Furthermore, at high rotation speeds, vortex configurations may be torn apart completely. Under those extreme circumstances, more exotic quantum phases may emerge, as seen in recent experimental and numerical studies identifying quantum Hall-like states [40, 41].
@@
Despite these advancements, most studies on rotating dipolar condensates have relied on mean-field descriptions based on the extended Gross-Pitaevskii equation (eGPE) [47]. While this approach captures essential features of dipolar Bose-Einstein condensates (BECs), it neglects crucial beyond-mean-field effects that become increasingly important in strongly correlated regimes [30]. In particular, the inclusion of the Lee-Huang-Yang (LHY) correction is often used to partially account for quantum fluctuations. The LHY term originates from the first-order correction to the ground-state energy of a weakly interacting Bose gas due to quantum fluctuations, as initially derived by Lee, Huang, and Yang for contact-interacting bosons in the dilute gas limit [48, 49]. This correction was later extended to dipolar gases, where it introduces an additional density-dependent term in the eGPE, modifying the effective equation of state of the condensate [50, 51].
@@
Despite great popularity of the eGPE, the phenomenological inclusion of the LHY correction is not fully self-consistent, as it is derived under the assumption of weak correlations and a homogeneous background density. In rapidly rotating condensates, where vortex lattices form and local density variations become significant, the LHY term derived in the homogeneous gas approximation may not accurately describe the system’s behavior [52, 53]. Moreover, in regimes where strong interactions lead to the formation of supersolid-like structures or cluster states, the validity of the LHY correction is further compromised, as it does not capture higher-order correlations, effective three-body interactions, or quantum depletion beyond Bogoliubov theory [19, 54]. This limitation underscores the need for beyond-mean-field approaches that can fully account for the many-body correlations in these systems.
@@
In addition to these theoretical limitations, the eGPE may also prove inadequate for describing certain experimental setups. Recent advances in experimental techniques have made it possible to study ultracold atomic gases consisting of only tens to hundreds of particles. Sophisticated trapping methods – ranging from optical pumping [55, 56] and optical tweezers [57] to geometric constrictions [58] and evaporative cooling [59] – have enabled precise control over small-scale quantum systems. In these small-scale systems, quantum fluctuations are enhanced due to the reduced particle number and stronger finite-size effects. The breakdown of the mean-field description in these cases suggests that fully beyond-mean-field approaches—such as quantum Monte Carlo methods, correlated many-body wavefunctions, or multiconfigurational time-dependent Hartree (MCTDH) methods—are necessary to capture the interplay of rotation, long-range interactions, and strong correlations.
@@
In this work, we employ the latter method to investigate the emergence of correlated phases in rotating dipolar condensates beyond the mean-field regime. By systematically varying interaction strength and rotation frequency for particle numbers from $N=5$ to $N=50$, we uncover a plethora of novel vortex structures, extended droplets, rotating cluster phases, and correlation-driven effects that cannot be captured by mean-field approximations. To quantify the deviations from mean-field theory, we compute an effective orbital occupation order parameter. Remarkably, we find that even in systems with a larger particle number beyond-mean-field effects become prominent when rotations are fast and/or interactions are strong.
@@ II System and protocol
The particles additionally feel a regularized dipole-dipole repulsion of strength $g_{d}$, described by
@@
with $\alpha=0.05$ a regularization parameter that arises in low-dimensional settings [60, 61, 62]. The regularization parameter induces a finite interaction in the limit $|\mathbf{r}_{j}-\mathbf{r}_{k}|\to 0$ and thus gives rise to a soft-core behavior that – in static settings – is known to induce density-modulated ground states [63, 64, 65, 66, 67]. In experiments, the interactions can be tuned via various mechanisms including confinement induced resonances [68, 69, 70, 71] and magnetic or electric field manipulation [72, 73, 74]. In particular, isotropic two-dimensional interactions can be achieved when the dipoles are confined to a plane and are all aligned perpendicularly to it. In our numerics, we adopt natural units corresponding to $\hbar=m=1$, and scale the Hamiltonian (13) by $\frac{\hbar^{2}}{mL^{2}}$, where $L$ is the natural length scale of the harmonic trap.
@@ III Methods
with the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles (MCTDH) [75, 76, 77, 78, 79, 80, 81]. The MCTDH approach assembles the many-body wave function $\Psi(\mathbf{r},t)$ as a dynamically optimized superposition of permanents. The permanents are in turn constructed from $M$ variationally determined single-particle functions, known as orbitals. The method self-consistently refines both the expansion coefficients and the orbital basis, facilitating ground-state computations via imaginary time relaxation or real-time evolution for dynamical studies. In this work, we focus exclusively on imaginary time propagation to ensure that the system reaches its lowest energy state. For numerical computations, we employ the MCTDH-X software suite [82, 83, 84, 85], which represents the state-of-the-art implementation of the MCTDH formalism. A detailed exposition of the method can be found in Appendix A.
@@
$\rho^{(1)}(\mathbf{r},\mathbf{r}^{\prime})$ =$\Psi\right|\hat{\Psi}^{\dagger}(\mathbf{r})\hat{\Psi}(\mathbf{r}^{\prime})\left|\Psi\right>.$ (8)
@@
$\rho(\mathbf{r})$ =$\Psi\right|\hat{\Psi}^{\dagger}(\mathbf{r})\hat{\Psi}(\mathbf{r})\left|\Psi\right>,$ (11)
@@ IV Results
Superfluid states (marked in white) are characterized by a condensed atomic cloud in a Thomas-Fermi shape [86] and an occupation of a single orbital near unity.
@@
As the rotation becomes stronger, it might become energetically favorable for the superfluid cloud to embed vortices in its core. Vortices are characterized by a phase winding of the superfluid order parameter around the vortex core and manifest in the density as depletion at the vortex center [87]. For sufficiently large numbers of vortices, they arrange into a triangular Abrikosov lattice [88, 89], which may be distorted due to finite-size effects and interactions [90, 91, 92]. In this work – to more precisely map the onset of vortex physics – we distinguish between single-vortex occurrences (marked in cyan in the phase diagrams) and multiple-vortex states (marked in green).
@@ IV.1 $N=50$ bosons
Figure 1: The different phases obtained from the simulation of $N=50$ dipolar-interacting bosons in a rotating potential as a function of interaction strength $g_{d}$ and rotation speed $\Omega$ (on a logarithmic scale). The different panels show results obtained for different number of orbitals: (a) $M=1$ orbitals, (b) $M=2$ orbitals, (c) $M=5$ orbitals.
@@
On top of this quantitative shift, though, using a larger number of orbitals also reveals qualitative differences in regimes of strong interactions, depicted in light blue and yellow in Fig. 1(c). The shape of these strongly interacting states depend on the rotational frequency. At slow rotations, the states appear as fragmented superfluids. These states display four-fold symmetric anisotropies in the density characteristic of angular roton softening [93], as depicted in Fig. 1(b), and can also coexist with single vortices in the center of the condensate, as shown in Fig. 1(d). The roton pattern appearing in the density is a precursor of a supersolid ordering that occurs at faster rotations [94, 95, 93]. In fact, in the strong interacting and fast rotating regime, we do observe the coexistence of clusters – arranged in a square lattice – with an underlying disordered vortex lattice. A prototypical configuration for this state is pictured in Fig. 2(f).
@@
Figure 5: The different phases obtained from the simulation of $N=20$ dipolar-interacting bosons in a rotating potential as a function of interaction strength $g_{d}$ and rotation speed $\Omega$ (on a logarithmic scale). The different panels show results obtained for different number of orbitals: (a) $M=1$ orbitals, (b) $M=2$ orbitals, (c) $M=5$ orbitals.
@@ IV.2 $N=20$ bosons
In fact, the density profiles depicted in Fig. 6 are fairly similar to the ones obtained for the $N=50$ calculations, with a few notable exceptions. The roton profile of the fragmented superfluids is stronger [6(b)], other fragmented states with a profile reminiscent of a radial roton [93] appear [6(c)], and the vortex lattices become more disordered [6(e)].
@@ IV.3 $N=10$ bosons
Figure 9: The different phases obtained from the simulation of $N=10$ dipolar-interacting bosons in a rotating potential as a function of interaction strength $g_{d}$ and rotation speed $\Omega$ (on a logarithmic scale). The different panels show results obtained for different number of orbitals: (a) $M=1$ orbitals, (b) $M=2$ orbitals, (c) $M=5$ orbitals. (d) $M=10$ orbitals.
@@
From the phase diagram, we immediately see that the mean-field calculations [Fig. 9(a)] give rise to a completely different picture than the calculations that account for quantum fluctuations [Fig. 9(b)-(d)]. This is expected, since $N=10$ are typically not enough particles to form stable superfluids beyond weakly-interacting cases [96]. In the mean-field phase diagram shown in Fig. 9(a), only superfluid states and vortices appear. These states have density profiles similar to the ones observed in systems with a larger number of particles, as illustrated in Fig. 10(a),(d). Accounting for correlations with $M>1$, though, gives rise to a progressively much richer landscape of states. The changes are particularly pronounced for fast rotations and strong interactions. Even with $M=2$ orbitals, a coexistence of clusters and vortices appears [yellow region, Fig. 9(b)]. Surprisingly, we find that the lattice formed by the clusters is not square, but triangular, as evidenced by the density in Fig. 10(e).
@@
Figure 10: The different states appearing for $N=20$ dipolar bosons in a rotating trap: (a) superfluid, (b) fragmented superfluid, (c) single-vortex threaded condensate, (d) multiple-vortex threaded condensate, (e) cluster lattice surrounded by vortices, (f) cluster lattice surrounded by aura, (g) single-particle clusters, (h) disrupted condensate (rotational speed too high). The circles in the upper-right corners have the same color code used in the phase diagram  9 and indicate which states the density represents.

### v2

@@ I Introduction
The recent progress in experiments with dipolar ultracold atoms [Griesmaier:2005, Lahaye:2008, Lu:2011, Aikawa:2012] and molecules [Ni:2008, Molony:2014, Park:2015, Rvachov:2017, Stevenson:2023, Bigagli:2024] has provided a platform for studying novel quantum phases emerging from the interplay between contact and long-range dipole-dipole interactions (DDI) [Lahaye:2009, Gadway:2016, Moses:2017]. These systems have enabled the realization of quantum droplets [Kartashov:2019, Zhang:2021] and supersolid phases [Lu:2015, Baillie:2018, Tanzi:2019, Chomaz:2019, Guo:2019, Natale:2019, Tanzi:2019-2, Tanzi:2021, Norcia:2021, Sohmen:2021, Biagioni:2022, Recati:2023, Chomaz:2023, Biagioni:2024], as well as the observation of dipolar quantum gases with metastable properties [Boettcher:2019, Sanchez-Baena:2023, Zhang:2024]. These interactions give rise to striking phenomena such as unconventional vortex clustering and dynamics [Easton:2023, Sabari:2024, Bland:2023, Tengstrand:2019], roton instabilities [Hufnagl:2011], and self-bound quantum droplets [Zhang:2021, Kartashov:2019].
@@
Beyond static properties, the interplay between dipolar interactions and external rotation introduces even richer physics. Recent studies have demonstrated that rotating dipolar gases exhibit complex vortex arrangements, turbulence, and dynamically stabilized cluster states [Kumar:2016, Bhowmik:2020, Fletcher:2021, Mukherjee:2022, Casotti:2024, Poli:2024, Chatterjee:2024]. A particularly intriguing aspect of rotating dipolar systems is their connection to astrophysical phenomena. In neutron stars, superfluid vortices exhibit long-range interactions that shape the internal structure and rotational properties of the star [Poli:2023, Bland:2024]. Understanding the interplay between rotation and dipolar interactions is therefore essential for advancing both laboratory-based quantum simulations and our knowledge of astrophysical superfluid systems.
@@
The presence of strong dipolar interactions modifies the well-known vortex lattice structures observed in non-dipolar BECs. In contrast to conventional Abrikosov lattices, dipolar condensates can exhibit striped vortex arrangements, fragmented clusters, or even phase-separated vortex distributions [Bland:2023, Sabari:2024]. Theoretical studies have predicted exotic vortex lattice phases, including triangular, square, and amorphous configurations, depending on the interaction strength and rotation frequency [Mukherjee:2022, Chatterjee:2024]. Furthermore, at high rotation speeds, vortex configurations may be torn apart completely. Under those extreme circumstances, more exotic quantum phases may emerge, as seen in recent experimental and numerical studies identifying quantum Hall-like states [Fletcher:2021, Mukherjee:2022].
@@
Despite these advancements, most studies on rotating dipolar condensates have relied on mean-field descriptions based on the extended Gross-Pitaevskii equation (eGPE) [Brachet:2012]. While this approach captures essential features of dipolar Bose-Einstein condensates (BECs), it neglects crucial beyond-mean-field effects that become increasingly important in strongly correlated regimes [Boettcher:2019]. In particular, the inclusion of the Lee-Huang-Yang (LHY) correction is often used to partially account for quantum fluctuations. The LHY term originates from the first-order correction to the ground-state energy of a weakly interacting Bose gas due to quantum fluctuations, as initially derived by Lee, Huang, and Yang for contact-interacting bosons in the dilute gas limit [Lee:1957-1, Lee:1957-2]. This correction was later extended to dipolar gases, where it introduces an additional density-dependent term in the eGPE, modifying the effective equation of state of the condensate [Lima:2011, Lima:2012].
@@
Despite great popularity of the eGPE, the phenomenological inclusion of the LHY correction is not fully self-consistent, as it is derived under the assumption of weak correlations and a homogeneous background density. In rapidly rotating condensates, where vortex lattices form and local density variations become significant, the LHY term derived in the homogeneous gas approximation may not accurately describe the system’s behavior [Bisset:2016, Waechtler:2016]. Moreover, in regimes where strong interactions lead to the formation of supersolid-like structures or cluster states, the validity of the LHY correction is further compromised, as it does not capture higher-order correlations, effective three-body interactions, or quantum depletion beyond Bogoliubov theory [Chomaz:2019, Hertkorn:2021]. This limitation underscores the need for beyond-mean-field approaches that can fully account for the many-body correlations in these systems.
@@
In addition to these theoretical limitations, the eGPE may also prove inadequate for describing certain experimental setups. Recent advances in experimental techniques have made it possible to study ultracold atomic gases consisting of only tens to hundreds of particles. Sophisticated trapping methods – ranging from optical pumping [Reetz-Lamour:2008-1, Reetz-Lamour:2008-2] and optical tweezers [Wang:2020] to geometric constrictions [Brantut:2012] and evaporative cooling [Zeiher:2021] – have enabled precise control over small-scale quantum systems. In these small-scale systems, quantum fluctuations are enhanced due to the reduced particle number and stronger finite-size effects. The breakdown of the mean-field description in these cases suggests that fully beyond-mean-field approaches—such as quantum Monte Carlo methods, correlated many-body wavefunctions, or multiconfigurational time-dependent Hartree (MCTDH) methods—are necessary to capture the interplay of rotation, long-range interactions, and strong correlations.
@@
In this work, employing the MCTDH method, we demonstrate that quantum correlations beyond mean-field theory play a crucial role in shaping the ground-state phase diagram of rotating dipolar condensates, even for relatively large particle numbers. By systematically varying interaction strength and rotation frequency for particle numbers from $N=5$ to $N=50$, we uncover a plethora of novel vortex structures, extended droplets, rotating cluster phases, and correlation-driven effects that cannot be captured by mean-field approximations. To quantify the deviations from mean-field theory, we compute an effective orbital occupation order parameter. Remarkably, we find that even in systems with a larger particle number beyond-mean-field effects become prominent when rotations are fast and/or interactions are strong.
@@ II System and protocol
The particles additionally feel a regularized dipole-dipole repulsion of strength $g_{d}$, effectively described by
@@
We note that – while a self-consistent dimensional reduction of the dipole-dipole interactions would lead to a normalization with Bessel functions [Sinha:2007, Ticknor:2009, Cremon:2010, Cai:2010, Deuretzbacher:2010, Yousefi:2015] – the simple regularization parameter $\alpha=0.05$ captures the short-range cutoff induced by transverse confinement equally well. Moreover, it ensures numerical stability without altering the qualitative phase behavior due to the universality of the dipole-dipole interactions [Hofmann:2021]. The regularization parameter induces a finite interaction in the limit $|\mathbf{r}_{j}-\mathbf{r}_{k}|\to 0$ and thus gives rise to a soft-core behavior that – in static settings – is known to induce density-modulated ground states [Cinti:2010, Cinti:2014, Biagioni:2018-thesis, Blakie:2024, Poli:2024-PRA]. In experiments, the interactions can be tuned via various mechanisms including confinement induced resonances [Tiesinga:1992, Tiesinga:2000, Bergeman:2003, Saeidian:2008] and magnetic or electric field manipulation [Tang:2018, Li:2021, Giovanazzi:2022]. In particular, isotropic two-dimensional interactions can be achieved when the dipoles are confined to a plane and are all aligned perpendicularly to it. In our numerics, we adopt natural units corresponding to $\hbar=m=1$, and scale the Hamiltonian (1) by $\frac{\hbar^{2}}{mL^{2}}$, where $L$ is the natural length scale of the harmonic trap.
@@ III Methods
with the MultiConfigurational Time-Dependent Hartree method for indistinguishable particles (MCTDH) [Streltsov:2006, Streltsov:2007, Alon:2007, Alon:2008, Lode:2012, Lode:2016, Fasshauer:2016]. The MCTDH approach assembles the many-body wave function $\Psi(\mathbf{r},t)$ as a dynamically optimized superposition of permanents. The permanents are in turn constructed from $M$ variationally determined single-particle functions, known as orbitals. The method self-consistently refines both the expansion coefficients and the orbital basis, facilitating ground-state computations via imaginary time relaxation or real-time evolution for dynamical studies. In this work, we focus exclusively on imaginary time propagation to ensure that the system reaches its lowest energy state. For numerical computations, we employ the MCTDH-X software suite [Lin:2020, Lode:2020, Molignini:2025-SciPost, MCTDHX], which represents the state-of-the-art implementation of the MCTDH formalism. A detailed exposition of the method can be found in Appendix A.
@@
$\rho^{(1)}(\mathbf{r},\mathbf{r}^{\prime})$ $=\left<\Psi\right|\hat{\Psi}^{\dagger}(\mathbf{r})\hat{\Psi}(\mathbf{r}^{\prime})\left|\Psi\right$>. (8)
@@
$\rho(\mathbf{r})$ $=\left<\Psi\right|\hat{\Psi}^{\dagger}(\mathbf{r})\hat{\Psi}(\mathbf{r})\left|\Psi\right$>, (11)
@@ IV Results
Superfluid states (marked in white) are characterized by a condensed atomic cloud in a Thomas-Fermi shape [Dalfovo:1999] and an occupation of a single orbital near unity.
@@
As the rotation becomes stronger, it might become energetically favorable for the superfluid cloud to embed vortices in its core. Vortices are characterized by a phase winding of the superfluid order parameter around the vortex core and manifest in the density as depletion at the vortex center [Fetter:2001]. For sufficiently large numbers of vortices, they arrange into a triangular Abrikosov lattice [Abrikosov:1957, Baym:2004], which may be distorted due to finite-size effects and interactions [Komineas:2007, Fetter:2008, Fetter:2009]. In this work – to more precisely map the onset of vortex physics – we distinguish between single-vortex occurrences (marked in cyan in the phase diagrams) and multiple-vortex states (marked in green).
@@ IV.1 $N=50$ bosons
Figure 1: The different phases obtained from the simulation of $N=50$ dipolar-interacting bosons in a rotating potential as a function of interaction strength $g_{d}$ and rotation speed $\Omega$ (on a logarithmic scale). The different panels show results obtained for different number of orbitals: (a) $M=1$ orbitals, (b) $M=2$ orbitals, (c) $M=5$ orbitals. The red letters in the phase diagram refer to the points in parameter space where the density distributions of Fig. 2 were probed.
@@
On top of this quantitative shift, though, using a larger number of orbitals also reveals qualitative differences in regimes of strong interactions, depicted in light blue and yellow in Fig. 1(c). The shape of these strongly interacting states depend on the rotational frequency. At slow rotations, the states appear as fragmented superfluids. These states display four-fold symmetric anisotropies in the density characteristic of angular roton softening [Schmidt:2021], as depicted in Fig. 2(b), and can also coexist with single vortices in the center of the condensate, as shown in Fig. 2(d). The roton pattern appearing in the density is a precursor of a supersolid ordering that occurs at faster rotations [Lahrz:2014, Alana:2021, Schmidt:2021]. In fact, in the strong interacting and fast rotating regime, we do observe the coexistence of clusters – arranged in a square lattice – with an underlying disordered vortex lattice. A prototypical configuration for this state is pictured in Fig. 2(f).
@@
Figure 5: The different phases obtained from the simulation of $N=20$ dipolar-interacting bosons in a rotating potential as a function of interaction strength $g_{d}$ and rotation speed $\Omega$ (on a logarithmic scale). The different panels show results obtained for different number of orbitals: (a) $M=1$ orbitals, (b) $M=2$ orbitals, (c) $M=5$ orbitals. The red letters in the phase diagram refer to the points in parameter space where the density distributions of Fig. 6 were probed.
@@ IV.2 $N=20$ bosons
In fact, the density profiles depicted in Fig. 6 are fairly similar to the ones obtained for the $N=50$ calculations, with a few notable exceptions. The roton profile of the fragmented superfluids is stronger [6(b)], other fragmented states with a profile reminiscent of a radial roton [Schmidt:2021] appear [6(c)], and the vortex lattices become more disordered [6(e)].
@@ IV.3 $N=10$ bosons
Figure 9: The different phases obtained from the simulation of $N=10$ dipolar-interacting bosons in a rotating potential as a function of interaction strength $g_{d}$ and rotation speed $\Omega$ (on a logarithmic scale). The different panels show results obtained for different number of orbitals: (a) $M=1$ orbitals, (b) $M=2$ orbitals, (c) $M=5$ orbitals. (d) $M=10$ orbitals. The red letters in the phase diagram refer to the points in parameter space where the density distributions of Fig. 10 were probed.
@@
From the phase diagram, we immediately see that the mean-field calculations [Fig. 9(a)] give rise to a completely different picture than the calculations that account for quantum fluctuations [Fig. 9(b)-(d)]. This is expected, since $N=10$ are typically not enough particles to form stable superfluids beyond weakly-interacting cases [Leggett:2001]. In the mean-field phase diagram shown in Fig. 9(a), only superfluid states and vortices appear. These states have density profiles similar to the ones observed in systems with a larger number of particles, as illustrated in Fig. 10(a),(d). Accounting for correlations with $M>1$, though, gives rise to a progressively much richer landscape of states. The changes are particularly pronounced for fast rotations and strong interactions. Even with $M=2$ orbitals, a coexistence of clusters and vortices appears [yellow region, Fig. 9(b)]. Surprisingly, we find that the lattice formed by the clusters is not square, but triangular, as evidenced by the density in Fig. 10(e).
@@
Figure 10: The different states appearing for $N=10$ dipolar bosons in a rotating trap: (a) superfluid, (b) fragmented superfluid, (c) single-vortex threaded condensate, (d) multiple-vortex threaded condensate, (e) cluster lattice surrounded by vortices, (f) cluster lattice surrounded by aura, (g) single-particle clusters, (h) disrupted condensate (rotational speed too high). The circles in the upper-right corners have the same color code used in the phase diagram  9 and indicate which states the density represents.